OpenStudy (anonymous):

Solve the limit: limit as h goes to infinity of [(1/h)] / [1+(1/h)] I know its one, but I can get it as 0/0 how can I proof its one? thanks

OpenStudy (anonymous):

limit as h goes to infinity of [(1/h)] / [1+(1/h)] = limit as h goes to infinity of [(1/h)] / [(h+1)/h)] = = limit as h goes to infinity of [(1/ [(h+1)] = 0

OpenStudy (anonymous):

crapp wrong one..its: Solve the limit: limit as h goes to infinity of [(1/h)] / [nl 1+(1/h)]

OpenStudy (anonymous):

i think this will just be zero